OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 5x - 3y = -3 2x - 6y = -6 A) -1 B) 0 C) 1 D) 2

OpenStudy (anonymous):

y=1 so C

OpenStudy (anonymous):

can i ask how you did this? @jcoolcat29

OpenStudy (anonymous):

2x - 6y = -6 2x = -6 + 6y x = -3 +3y Plug that answer into the other equation 5(-3 +3y) -3y = -3 -15 + 15y -3y = -3 12y = 12 y = 1 Does that make sense

OpenStudy (anonymous):

oh yeah yeah yeah ok i get that @jcoolcat29

OpenStudy (anonymous):

or i believe you can do it this way, if this helps any....